# Galaxy-scale lens search in the PEARLS NEP TDF and CEERS JWST fields

### v1

@@ 4 Results of the lens search
PEARLS172312+655014 $7.27_{-0.41}^{+0.01}$ $11.54_{-0.05}^{+0.06}$ $11.31_{-0.17}^{+0.17}$ $0.58_{-0.24}^{+0.39}$
PEARLS172239+655143 $7.67_{-0.38}^{+0.01}$ $11.61_{-0.11}^{+0.10}$ $11.11_{-0.08}^{+0.07}$ $0.31_{-0.11}^{+0.16}$
PEARLS172244+655623 $4.68_{-0.51}^{+0.01}$ $11.23_{-0.14}^{+0.17}$ $10.59_{-0.12}^{+0.12}$ $0.23_{-0.11}^{+0.19}$
CEERS142032+525823 $4.95_{-0.34}^{+0.01}$ $11.34_{-0.07}^{+0.06}$ $10.45_{-0.20}^{+0.12}$ $0.13_{-0.06}^{+0.07}$
@@
Figure 4: Comparison between the Einstein radius vs lens redshift distribution in our sample of lenses (left panel, orange and purple points), in COWLS (central panel, markers are mapped to the confidence score of each lens, S06 being the lowest and M25 the highest) and SL2S (right panel, red crosses). On top of each panel is reported the telescope used to conduct the relative lens search. The contours represent the iso-probability levels predicted using the galess analytical model (Ferrami & Wyithe 2024), including a cut in velocity dispersion ($\sigma_{\rm{SIE}}<210$ km/s) in the central panel. The shaded dashed contours in the right panel represent the expected distribution in SL2S if we assume the cut in velocity dispersion required to reproduce the distribution of COWLS.
@@ 5 Discussion and conclusions
We performed a search for galaxy-scale strong lenses in 144 arcmin2 of JWST NIRCam data, combining the PEARLS NEP TDF and CEERS fields, each containing observations in 9 photometric filters down to mag 28.5. We first visually inspected each NIRCam pointing, isolating 16 lens candidates, which we subsequentially tested for having different redshift between deflector and source with SED modelling, and for having image configurations compatible with a SIE mass distribution plus external shear. After applying these restrictions, we obtain a sample of 4 lenses, corresponding to a density of $100\pm 30$ lenses per deg2. We study the redshift and angular size distributions of this sample, comparing it with past searches and theoretical expectations. We confirm that a large fraction of the lenses discoverable by JWST ($\sim$ 30%) will be above redshift $z_{\rm{lens}}>1$ and have magnified sources with $z_{\rm{source}}>3$, such as PEARLS172239+655143 or the COSMOS-Web Ring (Mercier et al. 2024; van Dokkum et al. 2024; Shuntov et al. 2025), and most Einstein radii smaller than 1". Given the results of this article, we expect $\sim 25$ of these high-z lenses in COWLS (Nightingale et al. 2025) or a single Pure Parallel program (equivalent to around 200 NIRCam pointings). Such a set of high redshift lenses can be used to constrain the evolution of the elliptical galaxies mass function at redshift range higher what is available now. A sample of lensing systems with high redshifts and small angular sizes can also be used to train neural network lens finders on a range of parameters that has not been explored so far.
@@ acknowledgments
G.F and J.S.B.W. acknowledge the support of the Australian Research Council Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), through project number CE170100013, and the We acknowledge support from the ERC Advanced Investigator Grant EPOCHS (788113), as well as two studentships from STFC. J.M.D. acknowledges support from project PID2022-138896NB-C51 (MCIU/AEI/MINECO/FEDER, UE) Ministerio de Ciencia, Investigación y Universidades. This work is based on observations made with the NASA/ESA Hubble Space Telescope (HST) and NASA/ESA/CSA James Webb Space Telescope (JWST) obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute (STScI), which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST, and NAS 5–26555 for HST. The observations used in this work are associated with JWST programme 2738 and raw images are available on the MAST archive. The authors thank all involved in the construction and operations of the telescope as well as those who designed and executed these observations, their number are too large to list here and without each of their continued efforts, such work would not be possible.
@@
This work used the following software packages: astropy (Astropy Collaboration et al., 2013), pypher (Boucaud, A. et al., 2016), galfit (Peng et al., 2010), EAZY (Brammer et al., 2008), BAGPIPES (Carnall et al., 2018), lenstronomy (Birrer & Amara, 2018; Birrer et al., 2021), galess (Ferrami & Wyithe, 2024), NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007), SciPy (Virtanen et al., 2020).
@@
JWST data from PEARLS programme ID 2738 is publicly available on the MAST portal and accessible at this link. Accompanying HST data from TREASUREHUNT is available from https://doi.org/10.17909/wv13-qc14. JWST data from CEERS is publicly available at https://ceers.github.io/dr07.html.
@@ Appendix A Initial sample of lens candidates
Figure 6 contains the RGB composite images of the initial sample of 16 candidate lenses obtained by visual inspection over the PEARLS NEP TDF and CEERS JWST fields. The high confidence candidates, are highlighted with an orange border. Their photometric redshift measurements, where available, are listed in Table 4. Table 4 also lists the criteria that excluded a candidate from the final list of high probability lenses. If either of the photometric redshift is missing, or the photo-z measurements are incompatible with a lensing geometry, the rejection criteria is flagged as ‘photo-z’, while if the simple lens mass (SIE + external shear) or source light (single Sèrsic profile) models employed are not able to reproduce the observed configuration the rejection criteria is flagged as ‘lens model’.
Figure 6: Initial sample of lens candidates in PEARLS NEP field and CEERS fields, shown as 10"$\times$10" cutouts centred on the candidate deflector. The systems highlighted with an orange border are the 4 high-confidence strong lenses with measured photometric redshifts and a lens model.
@@
PEARLS172312+655014 17:23:11.93 +65:50:14.32 $0.3741_{-0.002}^{+0.002}$ $2.35_{-0.15}^{+0.05}$ (a)
PEARLS172239+655143 17:22:38.98 +65:51:43.16 $1.25_{-0.25}^{+0.25}$ $3.60_{-0.10}^{+0.10}$ (b)
PEARLS172244+655623 17:22:43.90 +65:56:22.97 $0.60_{-0.11}^{+0.09}$ $1.45_{-0.21}^{+0.14}$
PEARLS172340+654936 17:23:39.60 +65:49:35.75 $\left[0.38_{-0.10}^{+0.08}\right]$ $\left[2.60_{-0.20}^{+0.20}\right]$ lens model
PEARLS172233+654754 17:22:33.16 +65:47:53.66 $\left[0.57_{-0.15}^{+0.15}\right]$ $\left[2.90_{-0.14}^{+0.14}\right]$ lens model (c)
PEARLS172253+654847 17:22:53.14 +65:48:47.42 $\left[0.55._{-0.1}^{+0.1}\right]$ $\left[0.65._{-0.1}^{+0.1}\right]$ photo-z (d)
PEARLS172242+654338 17:22:42.68 +65:43:38.02 $\left[0.63._{-0.1}^{+0.1}\right]$ both (e)
PEARLS172245+655442 17:22:45.26 +65:54:41.96 $\left[0.98._{-0.1}^{+0.1}\right]$ both (e)
CEERS141956+525740 14:19:56.24 +52:57:40.80 - - both
CEERS141957+525926 14:19:57.30 +52:59:26.22 - - both
CEERS141933+525301 14:19:33.33 +52:53:00.90 $\left[1.13_{-0.17}^{+0.19}\right]$ $\left[1.51_{-0.13}^{+0.15}\right]$ lens model (f)
CEERS141857+524542 14:18:56.75 +52:45:42.45 - - photo-z (g)
CEERS141856+524528 14:18:55.60 +52:45:27.80 $\left[1.55_{-0.18}^{+0.18}\right]$ $\left[1.61_{-0.18}^{+0.18}\right]$ photo-z (d)
CEERS141938+525126 14:19:37.53 +52:51:25.97 - - lens model
CEERS142032+525823 14:20:31.80 +52:58:22.90 $0.380_{-0.005}^{+0.005}$ $0.99_{-0.05}^{+0.05}$ (h)
CEERS142021+525543 14:20:21.48 +52:55:43.12 $\left[0.73_{-0.24}^{+0.03}\right]$ $\left[4.21_{-0.21}^{+0.14}\right]$ lens model
Notes. From left to right, the columns indicate the lens name, right ascension and declination, photometric redshifts of the deflector and of the background source, and the criteria that excluded the system from the high-confidence sample (either photo-z or lens modeling). The systems without both redshift measurements or with measurements listed in square brackets are not included in the final sample of high-confidence lens candidates.

### v2

@@ 4 Results of the lens search
PEARLS J172311.9+655014 $7.27_{-0.41}^{+0.01}$ $1.54_{-0.05}^{+0.06} 11$$1.31_{-0.17}^{+0.17}$ $0.58_{-0.24}^{+0.39}$
PEARLS J172238.9+655143 $7.67_{-0.38}^{+0.01}$ $61_{-0.11}^{+0.10} 11.11$$11_{-0.08}^{+0.07}$ $0.31_{-0.11}^{+0.16}$
PEARLS J172243.9+655622 $4.68_{-0.51}^{+0.01}$ $11.23_{-0.14}^{+0.17}$ $10.59_{-0.12}^{+0.12}$ $0.23_{-0.11}^{+0.19}$
PEARLS J172339.6+654936 $7.94_{-1.49}^{+0.01}$ $1.61_{-0.13}^{+0.09} 11$$1.37_{-0.15}^{+0.18}$ $0.57_{-0.24}^{+0.60}$
CEERS J142031.8+525822 $4.95_{-0.34}^{+0.01}$ $11.34_{-0.07}^{+0.06}$ $10.45_{-0.20}^{+0.12}$ $0.13_{-0.06}^{+0.07}$
@@
Figure 5: Comparison between the Einstein radius vs lens redshift distribution in our sample of lenses (left panel, orange and purple points), in COWLS (central panel, markers are mapped to the confidence score of each lens, S06 being the lowest and M25 the highest) and SL2S (right panel, red crosses). On top of each panel is reported the telescope used to conduct the relative lens search. The contours represent the iso-probability levels predicted using the galess analytical model (Ferrami and Wyithe 2024), including a cut in velocity dispersion ($\sigma_{\rm{SIE}}<210$ km/s) in the central panel. The shaded dashed contours in the right panel represent the expected distribution in SL2S if we assume the cut in velocity dispersion required to reproduce the distribution of COWLS.
@@ 5 Discussion and conclusions
We performed a search for galaxy-scale strong lenses in 144 arcmin2 of JWST NIRCam data, combining the PEARLS NEP TDF and CEERS fields, each containing observations in 9 photometric filters down to mag 28.5. We first visually inspected each NIRCam pointing, isolating 16 lens candidates, which we subsequentially tested for having different redshift between deflector and source with SED modelling. Using a annular mask on each cutout, we fitted the lens and source light with Sérsic profiles and attempted to reproduced the observed image configurations with a SIE mass distribution plus external shear. To test whether the central portion of the mask and the light profile had any impact on our recovered sample, we remodelled the sample using a Multi-Gaussian Expansion to model the lens and source light down to the centre of the deflector. This technique lead to a lensing model of two additional systems compared to the previous step, performing particularly well on lenses with small angular size or faint lensed images compared to the lens light. Combining the requirements of a lens model and photo-zs compatible with lensing geometry, we obtained a sample of 5 high-confidence lenses, corresponding to a density of $125\pm 37$ lenses per deg2. We study the redshift and angular size distributions of this sample, comparing it with past searches and theoretical expectations. We confirm that a large fraction of the lenses discoverable by JWST ($\sim$ 30%) will be above redshift $z_{\rm{lens}}>1$ and have magnified sources with $z_{\rm{source}}>3$, such as PEARLS J172238.9+655143 or the COSMOS-Web Ring (Mercier et al. 2024; van Dokkum et al. 2024; Shuntov et al. 2025), and most Einstein radii smaller than 1". Given the results of this article, we expect $\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}30}$ of these high-z lenses in COWLS (Nightingale et al. 2025) or a single Pure Parallel program (equivalent to around 200 NIRCam pointings). Such a set of high redshift lenses can be used to constrain the evolution of the elliptical galaxies mass function at redshift range higher what is available now. A sample of lensing systems with high redshifts and small angular sizes can also be used to train neural network lens finders on a range of parameters that has not been explored so far.
@@ acknowledgments
We thank the referee, James W. Nightingale, for useful comments that improved the presentation. G.F and J.S.B.W. acknowledge the support of the Australian Research Council Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), through project number CE170100013, and the We acknowledge support from the ERC Advanced Investigator Grant EPOCHS (788113), as well as two studentships from STFC. J.M.D. acknowledges support from project PID2022-138896NB-C51 (MCIU/AEI/MINECO/FEDER, UE) Ministerio de Ciencia, Investigación y Universidades. This work is based on observations made with the NASA/ESA Hubble Space Telescope (HST) and NASA/ESA/CSA James Webb Space Telescope (JWST) obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute (STScI), which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST, and NAS 5–26555 for HST. The observations used in this work are associated with JWST programme 2738 and raw images are available on the MAST archive. The authors thank all involved in the construction and operations of the telescope as well as those who designed and executed these observations, their number are too large to list here and without each of their continued efforts, such work would not be possible.
@@
This work used the following software packages: astropy (Astropy Collaboration et al., 2013), pypher (Boucaud, A. et al., 2016), galfit (Peng et al., 2010), EAZY (Brammer et al., 2008), BAGPIPES (Carnall et al., 2018), lenstronomy (Birrer and Amara, 2018; Birrer et al., 2021), pyautolens (Nightingale and Dye, 2015; Nightingale et al., 2018, 2021), galess (Ferrami and Wyithe, 2024), NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007), SciPy (Virtanen et al., 2020).
@@
JWST data from PEARLS programme ID 2738 is publicly available on the MAST portal and accessible at this link. Accompanying HST data from TREASUREHUNT is available from https://doi.org/10.17909/wv13-qc14. JWST data from CEERS is publicly available at https://ceers.github.io/dr07.html. The FITS cutouts for all wavebands, including RMS noise maps and PSFs for all 16 candidates are made available at https://zenodo.org/records/15846915.
@@ Appendix A Initial sample of lens candidates
Figure 7 contains the RGB composite images of the initial sample of 16 candidate lenses obtained by visual inspection over the PEARLS NEP TDF and CEERS JWST fields. The high confidence candidates, are highlighted with an orange border. Their photometric redshift measurements, where available, are listed in Table 4. Table 4 also lists the criteria that excluded a candidate from the final list of high probability lenses. If either of the photometric redshift is missing, or the photo-z measurements are incompatible with a lensing geometry, the rejection criteria is flagged as ‘photo-z’, while if the simple lens mass (SIE + external shear) or source light (single Sèrsic profile) models employed are not able to reproduce the observed configuration the rejection criteria is flagged as ‘lens model’. After subtracting the lens light with a MGE, the systems CEERS J141856.7+524542, CEERS J141956.2+525740, PEARLS J172233.1+654753 appear as the best candidates for follow-up confirmation out of the 11 non-confirmed lensing systems.
Figure 7: Initial sample of lens candidates in PEARLS NEP field and CEERS fields, shown as $\times10"$ cutouts centred on the candidate deflector. The systems highlighted with an orange border are the 5 high-confidence strong lenses with measured photometric redshifts and a lens model.
@@
PEARLS J172311.9+655014 17:23:11.93 +65:50:14.32 $0.3741_{-0.002}^{+0.002}$ $2.35_{-0.15}^{+0.05}$ - (a)
PEARLS J172238.9+655143 17:22:38.98 +65:51:43.16 $1.25_{-0.25}^{+0.25}$ $3.60_{-0.10}^{+0.10}$ - (b)
PEARLS J172243.9+655622 17:22:43.90 +65:56:22.97 $0.60_{-0.11}^{+0.09}$ $1.45_{-0.21}^{+0.14}$ -
PEARLS J172339.6+654936 17:23:39.60 +65:49:35.75 $0.38_{-0.10}^{+0.08}$ $2.60_{-0.20}^{+0.20}$ -
PEARLS J172233.1+654753 17:22:33.16 +65:47:53.66 $\left[0.57_{-0.15}^{+0.15}\right]$ $\left[2.90_{-0.14}^{+0.14}\right]$ lens model (c)
PEARLS J172253.1+654847 17:22:53.14 +65:48:47.42 $\left[0.55._{-0.1}^{+0.1}\right]$ $\left[0.65._{-0.1}^{+0.1}\right]$ photo-z (d)
PEARLS J172242.6+654338 17:22:42.68 +65:43:38.02 $\left[0.63._{-0.1}^{+0.1}\right]$ - both (e)
PEARLS J172245.2+655441 17:22:45.26 +65:54:41.96 $\left[0.98._{-0.1}^{+0.1}\right]$ - both (e)
CEERS J141956.2+525740 14:19:56.24 +52:57:40.80 - - both
CEERS J141957.3+525926 14:19:57.30 +52:59:26.22 - - both
CEERS J141933.3+525300 14:19:33.33 +52:53:00.90 $\left[1.13_{-0.17}^{+0.19}\right]$ $\left[1.51_{-0.13}^{+0.15}\right]$ lens model (f)
CEERS J141856.7+524542 14:18:56.75 +52:45:42.45 - - photo-z (g)
CEERS J141855.6+524527 14:18:55.60 +52:45:27.80 $\left[1.55_{-0.18}^{+0.18}\right]$ $\left[1.61_{-0.18}^{+0.18}\right]$ photo-z (d)
CEERS J141937.5+525125 14:19:37.53 +52:51:25.97 - - lens model
CEERS J142031.8+525822 14:20:31.80 +52:58:22.90 $0.380_{-0.005}^{+0.005}$ $0.99_{-0.05}^{+0.05}$ - (h)
CEERS J142021.4+525543 14:20:21.48 +52:55:43.12 $\left[0.73_{-0.24}^{+0.03}\right]$ $\left[4.21_{-0.21}^{+0.14}\right]$ lens model
Notes. From left to right, the columns indicate the lens name, right ascension and declination, photometric redshifts of the deflector and of the background source, and the criteria that excluded the system from the high-confidence sample (either photo-z or lens modeling). The systems without both redshift measurements or with measurements listed in square brackets are not included in the final sample of high-confidence lens candidates. PEARLS J172311.9+655014 and PEARLS J172238.9+655143 are identified as radio sources in Hyun et al. 2023 with index 382 and 182, respectively.